OpenStudy (anonymous):

What is the Area of a circle with a radius of 5. Leave pi in your answer. please give me the answer dont explain because I barly have anytime to turn it in I will give you a medal and fan you, Thanks!!!

OpenStudy (anonymous):

@whpalmer4

OpenStudy (noseboy908):

Just answer me this: what's the formula for the area of a circle?

OpenStudy (whpalmer4):

What is the formula for the area of a circle?

OpenStudy (whpalmer4):

\[A = \pi r^2\]if you know the radius, or \[A = \frac{1}{4}\pi d^2\]if you know the diameter

OpenStudy (noseboy908):

I would typically wait for you to respond, but as you indicate that you're on a time crunch, I'll tell you it: "A= (pie)x(r^2)